40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Feb 3 15:57:52 2014, Steve, Update, Green Locking, ETMX green power ETMXgreen.jpgXgreenOven36C.jpg
    Reply  Tue Feb 4 01:02:03 2014, Koji, Update, Green Locking, ETMX green power Screenshot-Untitled_Window.png
    Reply  Tue Jan 19 17:27:58 2016, gautam, Update, Green Locking, AUX X power investigations X_END_POWER_BUDGET.JPG
Message ID: 11936     Entry time: Tue Jan 19 17:27:58 2016     In reply to: 9592
Author: gautam 
Type: Update 
Category: Green Locking 
Subject: AUX X power investigations 

Last week, Eric and I noticed that the green transmission levels at the PSL table seem much lower now than they did a month or two ago. To investigate this, I attempted to reproduce a power budget for the X endtable setup - see the attached figure (IR powers measured with calorimeter, green powers measured with Ophir power meter). A summary of my observations:

  • The measurements were all made at an AUX-X laser diode current of 1.90A, and laser crystal temperature of 47.41 degrees. The current was chosen on the basis of the AUX-X frequency noise investigations. The temperature was chosen as this is the middle of three end-laser temperatures at wich a beat-note can be found now. Why should this temperature have changed by almost 5 degrees from the value reported here? I checked on the PSL laser controller that the PSL temperature is 33.43 degrees. Turning up the diode current to 2A does not change the situation significantly. Also, on the Innolight datasheet, the tuning geometry graphs' X-axes only runs to 45 degrees. Not sure of what to make of this. I tried looking at the trend of the offset to the slow temperature servo to see if there has been some sort of long-term drift, but was unable to do so...
  • The IR power from the laser seems to have halved, compared to the value in Feb 2014. Is this normal deterioration over two years? Changing the laser diode current to 2A and the laser crystal temperature to ~42 degrees (the conditions under which the Feb 2014 measurements were taken) do not alter these numbers radically.
  • The green power seems to have become 1/4 its value in Feb 2014, which seems to be consistent with the fact that the IR power has halved.

It is worth noting that two years ago, the IR power from the AUX-Y laser was ~280 mW, so we should still be getting "enough" green power for ALS?

 

Attachment 1: X_END_POWER_BUDGET.JPG  432 kB  | Show | Show all
ELOG V3.1.3-